# Turning Sand to Gold: Recycling Data to Bridge On-Policy and Off-Policy Learning via Causal Bound

### v1

@@ C.5 MuJoCo Five Environments Benchmark
Table 11: Training rewards comparison between PPO agent using the additional SUFT OPE term and the baseline agent without it across 5 MuJoCo environments. SUFT surpasses the baseline agent in 4 out of 5 environments, with statistically significant gains in 2 environments. Both agents are using ${L_{2}}$ loss.
@@
Figure 15: Learning curves comparison between PPO agent using the additional SUFT OPE term and the baseline agent without it across 5 MuJoCo environments. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using ${L_{2}}$ loss.
@@ 8. Answer: [Yes]
  9. Justification: In Section 4, we describe the assumptions of our work, and in particular 4.1, which does not apply to the $L_{2}$ loss. In our experiments (Section 5), we evaluate our method on $L_{2}$ loss.
@@ 13. Answer: [Yes]
  14. Justification: In Section 4, we detailed the relevant assumptions for all theoretical results. We provide all proofs in Appendix B.
@@ 18. Answer: [Yes]
  19. Justification: We use open-source environments and an RL library as detailed in the Implementation Details paragraph in Section 5. Furthermore, we fully detailed all the experimental information in the Training Details paragraph in Section 5. Throughout Section 4 and Appendix B, we thoroughly detail how to implement our method. We also provide all the evaluation formulas in Appendix C. In addition, we also provide all of the code as supplementary material.
@@ 23. Answer: [Yes]
  24. Justification: We provide the code as supplementary material and fully detail the implementation and training details in Section 5. We also provide all the evaluation formulas in Appendix C.
@@ 28. Answer: [Yes]
  29. Justification: We fully detailed all the experimental information in the Training Details and Implementation Details paragraphs in Section 5.
@@ 38. Answer: [Yes]
  39. Justification: We give details on the computational time and resources in Appendix C.
@@ 58. Answer: [Yes]
  59. Justification: We cite all creators whose code we used in our experiments in Section 5.

### v2

@@ D.7 MuJoCo Five Environments Benchmark
Table 12: Training rewards comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 5 MuJoCo environments. SUFT surpasses the baseline agent in 4 out of 5 environments, with statistically significant gains in 2 environments. Both agents are using $L_{2}$} loss.
@@
Figure 16: Learning curves comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 5 MuJoCo environments. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{2}$} loss.
@@ 8. Answer: [Yes]
  9. Justification: In Section˜4, we describe the assumptions of our work, and in particular ˜4.1, which does not apply to the $L_{2}$ loss. In our experiments (Section˜5), we evaluate our method on $L_{2}$ loss.
@@ 13. Answer: [Yes]
  14. Justification: In Section˜4, we detailed the relevant assumptions for all theoretical results. We provide all proofs in Appendix˜B.
@@ 18. Answer: [Yes]
  19. Justification: We use open-source environments and an RL library as detailed in the Implementation Details paragraph in Section˜5. Furthermore, we fully detailed all the experimental information in the Training Details paragraph in Section˜5. Throughout Section˜4 and Appendix˜B, we thoroughly detail how to implement our method. We also provide all the evaluation formulas in Appendix˜D. In addition, we also provide all of the code as supplementary material.
@@ 23. Answer: [Yes]
  24. Justification: We provide the code as supplementary material and fully detail the implementation and training details in Section˜5. We also provide all the evaluation formulas in Appendix˜D.
@@ 28. Answer: [Yes]
  29. Justification: We fully detailed all the experimental information in the Training Details and Implementation Details paragraphs in Section˜5.
@@ 38. Answer: [Yes]
  39. Justification: We give details on the computational time and resources in Appendix˜D.
@@ 58. Answer: [Yes]
  59. Justification: We cite all creators whose code we used in our experiments in Section˜5.